1.       In 200A, the company’s sales was P500,000. Its fixed costs amounts to P100,000 per year. In 200B, sales was higher, while profit was P30,000 higher than the 200A figures.

 

For 200C, the company expects to have sales that is twice as much as the 200A sales. The expected increase in production to meet the sales demand in 200C will not require the company exceed its normal capacity.

 

Required:

a.       What is the company’s contribution margin ratio?

b.      How much profit does the company expect to earn in 200C?

c.       What is the company’s break-even point in units?

 


1
Expert's answer
2021-06-17T13:01:06-0400

Solution:

a.). Contribution margin ratio = "30\%"


Contribution margin = "\frac{Total\; sales - variable \; costs}{Total \; sales}"

200A = 500,000

200B = 500,000 + 120"\%" = 600,000

200C = 500,000 "\times" 2 = 1,000,000

Total Sales = 500,000 + 600,000 + 1,000,000 = 2,100,000

Variable costs = 1,470,000


Contribution margin = "\frac{2,100,000 - 1,470,000}{2,100,000} = 0.3\times100\% = 30\%"

 

b.). The company expects to earn 200,000 in 200C

Net income = Sales – variable costs – fixed costs

Net income = 1,000,000 – 700,000 – 100,000 = 200,000

Net income in 200C = 200,000

 

c.). The company’s break-even point in units cannot be determined using the given information